# A Cosmic Miracle: A Remarkably Luminous Galaxy at Confirmed with JWST

### v1

@@ Abstract
JWST has revealed a stunning population of bright galaxies at surprisingly early epochs, , where few such sources were expected. Here we present the most distant example of this class yet â MoM-z14, a luminous () source in the COSMOS legacy field at that expands the observational frontier to a mere 280 million years after the Big Bang. The redshift is confirmed with NIRSpec/prism spectroscopy through a sharp Lyman- break and detections of five rest-UV emission lines. The number density of bright sources implied by our âMirage or Miracleâ survey spanning arcmin2 is larger () than pre-JWST consensus models. The high EWs of UV lines (Ă ) signal a rising star-formation history, with a increase in the last 5 Myr (). The source is extremely compact (circularized pc), and yet resolved, suggesting an AGN is not the dominant source of light. The steep UV slope () implies negligible dust attenuation and a young stellar population. The absence of a strong damping wing may indicate that the immediate surroundings of MoM-z14 are partially ionized at a redshift where virtually every reionization model predicts a neutral fraction. The nitrogen emission and highly super-solar [N/C] hint at an abundance pattern similar to local globular clusters that may have once hosted luminous supermassive stars. Since this abundance pattern is also common among the most ancient stars born in the Milky Way, we may be directly witnessing the formation of such stars in dense clusters, connecting galaxy evolution across the entire sweep of cosmic time.
UTF8gbsn
@@ 1 Introduction
The quest to observe the earliest galaxies in the universe has been at the heart of observational cosmology since Hubbleâs discovery of the expanding universe a century ago. JWST was designed with this precise goal in mind: to push the observational frontier to cosmic dawn, revealing the first luminous objects that formed after the Big Bang (e.g., Gardner et al., 2006, 2023). Prior to JWSTâs launch, theoretical models predicted that detecting bright galaxies beyond redshift , past the Hubble Space Telescopeâs reach, would be extraordinarily challenging (e.g., Mason et al., 2015; Tacchella et al., 2018; Williams et al., 2018). For example, the sum total of deep and wide surveys planned for the first year of JWST observations (e.g., Treu et al., 2022; Eisenstein et al., 2023a; Finkelstein et al., 2025) were generally expected to yield only a handful of relatively faint ( mag AB) sources beyond that would take tens of hours of spectroscopy to confirm (e.g., Dayal et al., 2017; Yung et al., 2019; Behroozi et al., 2019). These predictions were built on hierarchical structure formation models, in which early galaxies were expected to be small, faint, and rare (e.g., White & Frenk, 1991; Bullock & Johnston, 2005; Springel et al., 2008). Broadly speaking, a baseline assumption was that the rate at which gas was transformed into stars (the star-formation efficiency; SFE), the modes of star-formation (e.g., âburstyâ vs. smooth), and the resulting types of stars and black holes at could be predicted based on what was observed at lower redshifts (e.g., DavĂ© et al., 2019; Vogelsberger et al., 2020; Kannan et al., 2022).
@@
Within weeks of the first science operations, JWSTâs images revealed an apparent abundance of bright galaxies at photometric redshifts, , challenging pre-JWST consensus models (e.g., Castellano et al., 2022; Naidu et al., 2022c; Finkelstein et al., 2022a; Atek et al., 2023; Harikane et al., 2023a). Over two years of photometric surveys continue to report this abundance of luminous sources at the highest redshifts (e.g., Casey et al., 2024; Donnan et al., 2024; Chemerynska et al., 2024; Adams et al., 2024; Kokorev et al., 2025a; Castellano et al., 2025a; PĂ©rez-GonzĂĄlez et al., 2025a). This unexpected population has electrified the community and raised fundamental questions about galaxy formation in the first Myrs. However, before we embark on revising the physics of the early Universe, systematic spectroscopic confirmation and characterization of these systems is necessary.
@@
Despite the deluge of candidates reported from photometric surveys, spectroscopic follow-up has been lagging, with only a relatively small number of sources confirmed at (e.g., Curtis-Lake et al., 2023; Bunker et al., 2023; Wang et al., 2023; Fujimoto et al., 2023; Harikane et al., 2024a; Hsiao et al., 2024; Napolitano et al., 2025). Concerningly, some of the most confidently selected sources have turned out to be low- interlopers, underscoring the critical need to systematically characterize the contamination fraction in photometric samples (e.g., Naidu et al., 2022c; Zavala et al., 2023; Donnan et al., 2023; Arrabal Haro et al., 2023; Harikane et al., 2025). At the bright end (), where GN-z11 and its peers reside, and at even higher redshifts where the challenge to models is the most acute in the face of the rapidly declining halo mass function, only a handful of sources have been confirmed to date (Castellano et al., 2024; Arrabal Haro et al., 2023; Carniani et al., 2024; Kokorev et al., 2025b). The most distant among these sources, JADES-GS-z14-0 (; Carniani et al. 2024, 2025a; Schouws et al. 2024) was found in a mere 10 arcmin2 survey implying a number density higher than the pre-JWST theoretical consensus (e.g., Tacchella et al., 2018; Robertson et al., 2024a; Whitler et al., 2025). Would this remarkable number density hold up if wider areas were systematically surveyed?
@@
The first spectra of luminous sources paint a stunningly rich portrait of the physics of these galaxies. Along many axes, there are few true analogs to these sources in the entire extragalactic Universe. For example, GNz11 (e.g., Bunker et al., 2023; Maiolino et al., 2024) and GLASS-z12/GHz2 (Naidu et al., 2022c; Castellano et al., 2022, 2024; Zavala et al., 2025; CalabrĂČ et al., 2024) display a remarkable chemical abundance pattern (e.g., super-solar [N/O]; e.g., Cameron et al. 2023a) that prior to JWST had been observed only in a handful of sources (Fosbury et al., 2003; PatrĂcio et al., 2016; Mingozzi et al., 2022). Might we be witnessing globular cluster formation in action (e.g., Senchyna et al., 2024)? Are these compact sources the highest redshift supermassive black holes (e.g., Maiolino et al., 2024)? Are there exotic stellar populations such as supermassive stars (SMS; ) at play (e.g., Charbonnel et al., 2023)? Large spectroscopic samples are needed to decide whether the extraordinary chemistry of these sources is ordinary (e.g., generic massive cluster formation; Belokurov & Kravtsov 2023).
@@
To address the need for spectroscopy at the cosmic frontier, we designed the âMirage or Miracleâ (MoM) JWST NIRSpec survey (GO-5224, PIs: Oesch & Naidu; Oesch et al., in prep.). We seek to test whether the extraordinary abundance of bright galaxies at the highest redshifts is a photometric mirage or a spectroscopic miracle, and whether their extraordinary physical conditions are peculiar or commonplace. MoM has systematically targeted a homogeneously selected sample of luminous galaxies across JWSTâs legacy wide-area fields (COSMOS, UDS) that were imaged with NIRCam+MIRI by an array of Cycle 1 and 2 programs, most prominently PRIMER (Donnan et al., 2024) and COSMOS-Web (Casey et al., 2023a) (see §3 for a full list of imaging programs in these fields utilized in the MoM target selection).
@@
Throughout this work, we adopt a flat CDM cosmology with parameters as per Planck Collaboration et al. (2018). We reference , the characteristic UV luminosity in Schechter function parametrizations of luminosity functions as per Bouwens et al. (2021) which corresponds to at . Magnitudes are in the AB system (e.g., Oke & Gunn, 1983). For summary statistics, we typically report medians with uncertainties on the median from bootstrapping (16th and 84th percentiles). We adopt the following values for solar abundances: , , 12+.
@@ 2 Data
MoMâs primary goal is to investigate the nature of luminous () photometric candidates. Each of the five NIRSpec pointings in this program â two in the UDS field and three in COSMOS â is designed around such candidates as the primary targets. Every pointing receives 4.4 hrs of integration in the prism mode. At this depth, the putative Ly break in our candidates may be distinguished from Balmer breaks as well as low- emission line galaxies. Further details of the MoM sample, survey design, and full survey yield will be described in Oesch et al., in prep.
@@ 2.1 Imaging, Photometry, and Selection of MoM-z14
To build the parent catalog for MoM targets, we combined all publicly available JWST imaging over the COSMOS field as of January 2025 arising from the following programs: COSMOS-Web (#1727; Casey et al. 2023b), Blue Jay (#1810; Belli et al. 2024), PRIMER (#1837; Donnan et al. 2024), JELS (#2321; Duncan et al. 2024), PANORAMIC (#2514; Williams et al. 2025), COSMOS-3D (#5893; Kakiichi et al. 2024), and SAPPHIRES (#6434; Sun et al. 2025). These images are publicly released on the DAWN JWST Archive (DJA)111https://dawn-cph.github.io/dja/index.html as v7.4 COSMOS mosaics. Details about the imaging data reduction using the grizli software (Brammer, 2018; Brammer et al., 2022) may be found in Valentino et al. (2023). PSF-matched photometric catalogs based on these images were produced following the procedure outlined in Weibel et al. (2024). Fluxes for the source of interest are listed in Table 1.
We fit photometric redshifts to identify potential targets using eazy (Brammer et al., 2008a, b) with the blue_sfhz template set featuring thirteen FSPS templates (Conroy et al., 2009, 2010; Conroy & Gunn, 2010b, a; Foreman-Mackey et al., 2014) and an empirical template based on the strong emission line source at studied in Carnall et al. (2023). We also include a template from Naidu et al. (2022a) fit to a dusty emission line galaxy that virtually all photo- routines place at with very high confidence (see discussion in Arrabal Haro et al. 2023; Donnan et al. 2023; Harikane et al. 2024a; Zavala et al. 2023).
@@ 2.2 Spectroscopy
MoM-z14 was targeted in one of the COSMOS MoM masks observed on 16th April, 2025. We reduced its spectrum using v0.9.4 of the msaexp software (Brammer, 2023), details of which can be found in Heintz et al. (2024b); de Graaff et al. (2024a). The reduction choices we employ are the exact same as the upcoming v4 NIRSpec release of the DAWN JWST Archive (Pollock et al., in prep.). A key development in this version compared to Heintz et al. (2024b); de Graaff et al. (2024a) is an updated flux calibration and an expansion of the nominal prism wavelength coverage up to m. The processed prism spectrum (4.4h exposure time) is shown in Fig. 1 along with the well-centered position of the source in its assigned MSA slitlets.
We perform two data validation checks that are shown in the Appendix. First, the shape and normalization of the prism spectrum is in excellent agreement with the NIRCam photometry (Fig. 11). Note that the photometry is not used in any step of the data reduction, and so this is an independent validation of the overall flux calibration. Next, we reduce the spectrum with a âglobalâ sky background where the sky spectrum is estimated from empty slitlets on the mask. This is in contrast to the fiducial spectrum based on the âlocalâ differencing at different nod positions. Given the high background in COSMOS owing to its proximity to the ecliptic (relative to, e.g., EGS, Abell 2744, GOODS-N), this is a potentially important systematic. Importantly, all the key features analyzed in this work (e.g., the sharp break, UV emission lines) are detected in both versions (see Fig. 12).
@@
 (rest-frame). FWHM for all lines is allowed to range up to 650 km s-1, with the width of all lines tied to each other â the FWHM is completely unconstrained in our fits. aQuintuplet comprising lines at 1747Ă
, 1749Ă
, 1750Ă
, 1752Ă
, 1754Ă
@@ 3.1.1 UV lines
We use UNITE, a custom NIRSpec emission line fitting package (Hviding et al., in prep.), to model the UV emission lines. This package self-consistently handles the translation between idealized models of the spectrum to realistic observed NIRSpec spectra by accounting for e.g., the wavelength-dependent resolution, LSF (idealized for a point-source, applicable to our relatively compact object) and calibration uncertainties following de Graaff et al. (2024b, c). We initialize Gaussians to represent the typical UV lines seen in spectra of luminous galaxies (e.g., Bunker et al., 2023; Maiolino et al., 2024; Castellano et al., 2024) with uniform priors on their FWHM up to 650 km s-1 and redshift between . The width and redshift are tied across all lines. Various multiplets and adjacent lines are blended at the prism resolution â we model these by splitting the lines but report combined fluxes by summing their posterior fluxes. The continuum is fit by considering a region of km s-1 around each line, and masking a central region of km s-1. Overlapping regions around adjacent emission lines are stitched together and treated as a contiguous region. The posteriors are sampled using numpyro (Phan et al., 2019) and are shown in Fig. 2 and reported in Table 2.2.
We recover five UV lines at (top-right panel, Fig. 2). Reassuringly, the reported lines correspond to peaks in the 2D spectrum, and are also seen in a version of the spectrum processed using the global sky background (Fig. 12). The strongest lines detected â forbidden and semi-forbidden transitions of nitrogen and Carbon â also happen to be the strongest UV lines seen in other comparably luminous galaxies (e.g., Bunker et al. 2023; Castellano et al. 2024; Carniani et al. 2024).
Owing to these UV lines, the inferred redshift is quite precise (), which is unique for a prism redshift at this early epoch (e.g., Curtis-Lake et al., 2023; Wang et al., 2023; Carniani et al., 2024). This will enable extremely efficient ALMA spectral scan follow-up for e.g., [Oâiii]88m and [Câii]158m as well as precise modeling of the damping wing. Typically, due to the paucity of detected lines at prism resolution, only Ly break redshifts with much wider posteriors have been possible (see bottom panels of Fig. 2). This is due to degeneracies between the location of the break, the neutral fraction of the IGM, and the presence of damped Ly absorption â e.g., for JADES-GS-z14-0 subsequently refined with ALMA to (Carniani et al., 2024; Schouws et al., 2024; Carniani et al., 2025a).

### v2

@@ Abstract
JWST has revealed a stunning population of bright galaxies at surprisingly early epochs, , where few such sources were expected. Here we present the most distant example of this class yet – MoM-z14, a luminous () source in the COSMOS legacy field at that expands the observational frontier to a mere 280 million years after the Big Bang. The redshift is confirmed with NIRSpec/prism spectroscopy through a sharp Lyman- break and detections of five rest-UV emission lines. The number density of bright sources implied by our “Mirage or Miracle” survey spanning arcmin2 is larger () than pre-JWST consensus models. The high EWs of UV lines (Å) signal a rising star-formation history, with a increase in the last 5 Myr (). The source is extremely compact (circularized pc), and yet elongated (), suggesting an AGN is not the dominant source of UV light. The steep UV slope () implies negligible dust attenuation and a young stellar population. The absence of a strong damping wing provides tentative evidence that the immediate surroundings of MoM-z14 may be partially ionized at a redshift where virtually every reionization model predicts a neutral fraction. The nitrogen emission and highly super-solar [N/C] hint at an abundance pattern similar to local globular clusters that may have once hosted luminous supermassive stars. Since this abundance pattern is also common among the most ancient stars born in the Milky Way, we may be directly witnessing the formation of such stars in dense clusters, connecting galaxy evolution across the entire sweep of cosmic time.
@@ 1 Introduction
The quest to observe the earliest galaxies in the universe has been at the heart of observational cosmology since Hubble’s discovery of the expanding universe a century ago. JWST was designed with this precise goal in mind: to push the observational frontier to cosmic dawn, revealing the first luminous objects that formed after the Big Bang (e.g., Gardner et al., 2006, 2023). Prior to JWST’s launch, theoretical models predicted that detecting bright galaxies beyond redshift , past the Hubble Space Telescope’s reach, would be extraordinarily challenging (e.g., Mason et al., 2015; Tacchella et al., 2018; Williams et al., 2018). For example, the sum total of deep and wide surveys planned for the first year of JWST observations (e.g., Treu et al., 2022; Eisenstein et al., 2023b; Finkelstein et al., 2025) were generally expected to yield only a handful of relatively faint ( mag AB) sources beyond that would take tens of hours of spectroscopy to confirm (e.g., Dayal et al., 2017; Yung et al., 2019; Behroozi et al., 2019). These predictions were built on hierarchical structure formation models, in which early galaxies were expected to be small, faint, and rare (e.g., White and Frenk, 1991; Bullock and Johnston, 2005; Springel et al., 2008). Broadly speaking, a baseline assumption was that the rate at which gas was transformed into stars (the star-formation efficiency; SFE), the modes of star-formation (e.g., “bursty” vs. smooth), and the resulting types of stars and black holes at could be predicted based on what was observed at lower redshifts (e.g., Davé et al., 2019; Vogelsberger et al., 2020; Kannan et al., 2022).
@@
Within weeks of the first science operations, JWST’s images revealed an apparent abundance of bright galaxies at photometric redshifts, , challenging pre-JWST consensus models (e.g., Castellano et al., 2022; Naidu et al., 2022c; Finkelstein et al., 2022b; Atek et al., 2023; Harikane et al., 2023a). Over two years of photometric surveys continue to report this abundance of luminous sources at the highest redshifts (e.g., Casey et al., 2024; Donnan et al., 2024; Chemerynska et al., 2024; Adams et al., 2024; Kokorev et al., 2025a; Castellano et al., 2025a; Pérez-González et al., 2025a). This unexpected population has electrified the community and raised fundamental questions about galaxy formation in the first Myrs. However, before we embark on revising the physics of the early Universe, systematic spectroscopic confirmation and characterization of these systems is necessary.
@@
Despite the deluge of candidates reported from photometric surveys, spectroscopic follow-up has been lagging, with only a relatively small number of sources confirmed at (e.g., Curtis-Lake et al., 2023; Bunker et al., 2023; Wang et al., 2023; Fujimoto et al., 2023; Harikane et al., 2024b; Hsiao et al., 2024; Napolitano et al., 2025). Concerningly, some of the most confidently selected sources have turned out to be low- interlopers, underscoring the critical need to systematically characterize the contamination fraction in photometric samples (e.g., Naidu et al., 2022c; Zavala et al., 2023; Donnan et al., 2023; Arrabal Haro et al., 2023; Harikane et al., 2025). At the bright end (), where GN-z11 and its peers reside, and at even higher redshifts where the challenge to models is the most acute in the face of the rapidly declining halo mass function, only a handful of sources have been confirmed to date (Castellano et al., 2024; Arrabal Haro et al., 2023; Carniani et al., 2024; Kokorev et al., 2025b). The most distant among these sources, JADES-GS-z14-0 (; Carniani et al. 2024, 2025a; Schouws et al. 2024) was found in a mere 10 arcmin2 survey implying a number density higher than the pre-JWST theoretical consensus (e.g., Tacchella et al., 2018; Robertson et al., 2024a; Whitler et al., 2025). Would this remarkable number density hold up if wider areas were systematically surveyed?
@@
The first spectra of luminous sources paint a stunningly rich portrait of the physics of these galaxies. Along many axes, there are few true analogs to these sources in the entire extragalactic Universe. For example, GNz11 (e.g., Bunker et al., 2023; Maiolino et al., 2024) and GLASS-z12/GHz2 (Naidu et al., 2022c; Castellano et al., 2022, 2024; Zavala et al., 2025; Calabrò et al., 2024) display a remarkable chemical abundance pattern (e.g., super-solar [N/O]; e.g., Cameron et al. 2023a) that prior to JWST had been observed only in a handful of sources (Fosbury et al., 2003; Patrício et al., 2016; Mingozzi et al., 2022). Might we be witnessing globular cluster formation in action (e.g., Senchyna et al., 2024)? Are these compact sources the highest redshift supermassive black holes (e.g., Maiolino et al., 2024)? Are there exotic stellar populations such as supermassive stars (SMS; ) at play (e.g., Charbonnel et al., 2023)? Large spectroscopic samples are needed to decide whether the extraordinary chemistry of these sources is ordinary (e.g., generic massive cluster formation; Belokurov and Kravtsov 2023).
@@
To address the need for spectroscopy at the cosmic frontier, we designed the “Mirage or Miracle” (MoM) JWST NIRSpec survey (GO-5224, PIs: Oesch & Naidu; Oesch et al., in prep.). We seek to test whether the extraordinary abundance of bright galaxies at the highest redshifts is a photometric mirage or a spectroscopic miracle, and whether their extraordinary physical conditions are peculiar or commonplace. MoM has systematically targeted a homogeneously selected sample of luminous galaxies across JWST’s legacy wide-area fields (COSMOS, UDS) that were imaged with NIRCam+MIRI by an array of Cycle 1 and 2 programs, most prominently PRIMER (Donnan et al., 2024) and COSMOS-Web (Casey et al., 2023a) (see §3 for a full list of imaging programs in these fields utilized in the MoM target selection).
@@
Throughout this work, we adopt a flat CDM cosmology with parameters as per Planck Collaboration et al. (2018). We reference , the characteristic UV luminosity in Schechter function parametrizations of luminosity functions as per Bouwens et al. (2021) which corresponds to at . Magnitudes are in the AB system (e.g., Oke and Gunn, 1983). For summary statistics, we typically report medians with uncertainties on the median from bootstrapping (16th and 84th percentiles). We adopt the following values for solar abundances: , , 12+.
@@ 2 Data
MoM’s primary goal is to investigate the nature of luminous () photometric candidates. Each of the five NIRSpec pointings in this program – two in the UDS field and three in COSMOS – is designed around such candidates as the primary targets. Every pointing receives 4.4 hrs of integration in the prism mode. At this depth, the putative Ly break in our candidates may be distinguished from Balmer breaks as well as low- emission line galaxies. Further details of the MoM sample, survey design, and full survey yield will be described in Oesch et al., in prep.
@@ 2.1 Imaging, Photometry, and Selection of MoM-z14
To build the parent catalog for MoM targets, we combined all publicly available JWST imaging over the COSMOS field as of January 2025 arising from the following programs: COSMOS-Web (#1727; Casey et al. 2023b), Blue Jay (#1810; Belli et al. 2024), PRIMER (#1837; Donnan et al. 2024), JELS (#2321; Duncan et al. 2024), PANORAMIC (#2514; Williams et al. 2025), COSMOS-3D (#5893; Kakiichi et al. 2024), and SAPPHIRES (#6434; Sun et al. 2025). These images are publicly released on the DAWN JWST Archive (DJA)111https://dawn-cph.github.io/dja/index.html as v7.4 COSMOS mosaics. Details about the imaging data reduction using the grizli software (Brammer, 2018; Brammer et al., 2022) may be found in Valentino et al. (2023). PSF-matched photometric catalogs based on these images were produced following the procedure outlined in Weibel et al. (2024). Fluxes for the source of interest are listed in Table 1. These are fluxes derived with an adopted aperture radius of that are then corrected to total following Weibel et al. (2024).
We fit photometric redshifts to identify potential targets using eazy (Brammer et al., 2008a, b) with the blue_sfhz template set featuring thirteen FSPS templates (Conroy et al., 2009, 2010; Conroy and Gunn, 2010b, a; Foreman-Mackey et al., 2014) and an empirical template based on the strong emission line source at studied in Carnall et al. (2023). We also include a template from Naidu et al. (2022b) fit to a dusty emission line galaxy that virtually all photo- routines place at with very high confidence (see discussion in Arrabal Haro et al. 2023; Donnan et al. 2023; Harikane et al. 2024b; Zavala et al. 2023).
@@ 2.2 Spectroscopy
MoM-z14 was targeted in one of the COSMOS MoM masks observed on 16th April, 2025. We reduced its spectrum using v0.9.4 of the msaexp software (Brammer, 2023), details of which can be found in Heintz et al. (2024a); de Graaff et al. (2024c). The reduction choices we employ are the exact same as the upcoming v4 NIRSpec release of the DAWN JWST Archive (Pollock et al., in prep.; Valentino et al. 2025). A key development in this version compared to Heintz et al. (2024a); de Graaff et al. (2024c) is an updated flux calibration and an expansion of the nominal prism wavelength coverage up to m. The processed prism spectrum (4.4h exposure time) is shown in Fig. 1 along with the well-centered position of the source in its assigned MSA slitlets.
We perform two data validation checks that are shown in the Appendix. First, the shape and normalization of the prism spectrum is in excellent agreement with the NIRCam photometry (Fig. 12). Note that the photometry is not used in any step of the data reduction, and so this is an independent validation of the overall flux calibration. Next, we reduce the spectrum with a “global” sky background where the sky spectrum is estimated from empty slitlets on the mask. This is in contrast to the fiducial spectrum based on the “local” differencing at different nod positions. Given the high background in COSMOS owing to its proximity to the ecliptic (relative to, e.g., EGS, Abell 2744, GOODS-N), this is a potentially important systematic. Importantly, all the key features analyzed in this work (e.g., the sharp break, UV emission lines) are detected in both versions (see Fig. 13).
@@ 3.1.1 UV lines
We use UNITE, a custom NIRSpec emission line fitting package (Hviding et al., 2025), to model the UV emission lines. This package self-consistently handles the translation between idealized models of the spectrum to realistic observed NIRSpec spectra by accounting for e.g., the wavelength-dependent resolution, LSF (idealized for a point-source, applicable to our relatively compact object) and calibration uncertainties following de Graaff et al. (2024b, a). We initialize Gaussians to represent the typical UV lines seen in spectra of luminous galaxies (e.g., Bunker et al., 2023; Maiolino et al., 2024; Castellano et al., 2024) with uniform priors on their FWHM up to 650 km s-1 and redshift between . The width and redshift are tied across all lines. By construction, with UNITE the fitted widths of the emission lines are consistent with the NIRSpec LSF. Various multiplets and adjacent lines are blended at the prism resolution – we model these by splitting the lines in a multi-Gaussian fit but report combined fluxes by summing their posterior fluxes. The continuum is fit by considering a region of km s-1 around each line, and masking a central region of km s-1. Overlapping regions around adjacent emission lines are stitched together and treated as a contiguous region. The posteriors are sampled using numpyro (Phan et al., 2019) and are shown in Fig. 2 and reported in Table 2.2.
We recover five UV lines at (top-right panel, Fig. 2). There are also hints of additional lines (e.g., Ne iv and C ii) albeit at lower significance () that will require deeper data to confirm. Reassuringly, the reported lines correspond to peaks in the 2D spectrum, and are also seen in a version of the spectrum processed using the global sky background (Fig. 13). The strongest lines detected – forbidden and semi-forbidden transitions of nitrogen and Carbon – also happen to be the strongest UV lines seen in other comparably luminous galaxies (e.g., Bunker et al. 2023; Castellano et al. 2024; Carniani et al. 2024).
Owing to these UV lines, the inferred redshift is quite precise (), which is unique for a prism redshift at this early epoch (e.g., Curtis-Lake et al., 2023; Wang et al., 2023; Carniani et al., 2024). This will enable extremely efficient ALMA spectral scan follow-up for e.g., [O iii]88m and [C ii]158m as well as precise modeling of the damping wing. Typically, due to the paucity of detected lines at prism resolution, only Ly break redshifts with much wider posteriors have been possible (see bottom panels of Fig. 2). This is due to degeneracies between the location of the break, the neutral fraction of the IGM, and the presence of damped Ly absorption – e.g., for JADES-GS-z14-0 subsequently refined with ALMA to (Carniani et al., 2024; Schouws et al., 2024; Carniani et al., 2025a).